OpenStudy (solomonzelman):

What do you think? I was doing a short project and needed a Ti84 calculator real quick. What I did was...

OpenStudy (solomonzelman):

I bought this calculator (in Target), and then USED IT, opening the stuff it was packed in, so that I can put it back in the same way as if I never opened it, and gave it back to the store as if I didn't need it. It was a very expensive calculator, 135 bucks...

OpenStudy (anonymous):

Yeah I had to buy one for high school haha (was $200 at the time), I actually don't even remember how to use it now. It's kind of ridiculous though, here they teach you how to graph on the calculator, rather than on paper...and they ask why foreign students are geniuses haha.

OpenStudy (solomonzelman):

I had to make a sinusoidal regression of a bunch of points, but having a computer I found a you-tube video with explanations.

OpenStudy (solomonzelman):

(with an explanation about how to use the calc)

OpenStudy (anonymous):

That's good for that reason I suppose. But, when I was in like grade 11, they were just like press this button here and insert that here, and you have a graph. They never explained equation of a straight line and etc, I had to learn that all myself when I first started University...

OpenStudy (anonymous):

That's why in calculus 50% of the class fails/ drops out, I blame the high school teachers haha.

OpenStudy (solomonzelman):

Yeah... I hate learning things without a teacher. Even the past year, when I took home school, it took about a month to figure out how to use the school's website.

OpenStudy (solomonzelman):

─ :(

OpenStudy (anonymous):

Yeah it sucks, but now I've learnt how important it is I guess lol, and I can say I do enjoy learning some of the things myself, but of course a good teacher would be much more helpful. Doing research yourself, I believe makes the information stick better, well for me, I'm not sure about the others.

OpenStudy (solomonzelman):

IT just depends, when I find what I needed I member it, but when others do it for me, then I don't understand it just like before.

OpenStudy (anonymous):

Mostly this happens because I don't practice enough lol, I skip a lot of the homework, yet I still manage. But as I mentioned before, I'm really enjoying learning now, in a way I have this thirst for knowledge that I didn't before, guess it hit me soon as I started University lol.

OpenStudy (anonymous):

Back to the topic, I think you could've used some online calculators rather than having to buy the ti84 now right? Wouldn't wolframalpha suffice?

OpenStudy (solomonzelman):

Wolfram has Ti84 calc? I tried to enter points for a regression into it, but it gave me just the plotted coordinates of the points.

OpenStudy (anonymous):

Mhm actually I'm not entirely sure, I haven't messed around with wolfram much, but I believe there is regression on it. Also, now days you can find those calculators (emulators) online *cough* google search *cough* haha.

OpenStudy (solomonzelman):

Yeah... I tried to google for it, but nothing efficient seemed to come up fast enough.

OpenStudy (anonymous):

Ah, well least it worked out for ya, did target accept the calculator back haha?

OpenStudy (solomonzelman):

Oh yeah:) I would probably be working somewhere and making up for the lost money if the target didn't accept it back

OpenStudy (anonymous):

Haha, nice! :D glad it all worked out.

OpenStudy (solomonzelman):

ty:)

thomaster (thomaster):

you know there are lots of online graphing calculators that work the same way right? :P https://www.desmos.com/calculator

OpenStudy (solomonzelman):

yes, I am actually a big fan of desmos , but when I just given the points, it's unable to calculate a sinusoidal regression.

thomaster (thomaster):

how about wolframalpha?

OpenStudy (amistre64):

hmm, a regression equation is simply the least squared errors; which resorts to partial derivatives\[f(A,\omega,\rho,B)=\sum(Asin(\omega(x_i-\rho))+B-y_i)^2\] \[f_A=\sum2sin(\omega(x_i-\rho))~(Asin(\omega(x_i-\rho))+B-y_i)\] \[f_\omega=\sum2A(x_i-\rho)~cos(\omega(x_i-\rho))~(Asin(\omega(x_i-\rho))+B-y_i)\] \[f_\rho=\sum-2A\omega~cos(\omega(x_i-\rho))~(Asin(\omega(x_i-\rho))+B-y_i)\] \[f_B=\sum2(Asin(\omega(x_i-\rho))+B-y_i)\] then setting each partial equal to zero, we develop a system of equations to work with.